# Field Matters: A lightweight LLM-enhanced Method for CTR Prediction

### v1

@@ A.3 Baselines
  • KAR (RecSys’24) [8]: KAR leverages the LLM as a open-word knowledge base to enhance the profile of users and items, and then pass the encoded user/item embeddings through a knowledge adaptation network to enhance traditional CTR models.
@@
  • LLM-CF (CIKM’24) [9]: LLM-CF uses LLMs to generate a base of chains of thought (CoT) for sampled instances, which are further retrieved to enhance CTR models. Consistent with the original paper, we sampled 10% of the instances as sampled data for fine-tuning and CoT data generation, and during downstream CTR prediction, we also retrieve the Top 4 samples closest in semantics from the CoT dataset as enhanced knowledge.
@@
  • CTRL (TORS’23) [7]: CTRL utilizes the LLM as the feature encoder for instances and uses a fine-grained contrastive learning scheme to align cross-modal knowledge. In our implementation, we used Llama3-8B [3] as the LLM to encode instances embeddings for fair comparison.
@@
  • EASE (CIKM’24) [10]: EASE trains a semantic adaptor network to align item embeddings with the LLM, and then use it for encoding during inference. Following the original paper, we inited the semantic adapter using BERT-Base [68] with 12 transformer blocks, and cross-attention layers are added to the 6-th and 12-th transformer blocks.
@@
For fair comparisons, unless otherwise specified, we consistently used the Llama3-8B [3] as the LLM in all the LLM-enhanced CTR methods.
Besides, we also reproduced two representative PLM-enhanced methods:
@@
  • CELA (arXiv’24) [12]: CELA first pretrains a PLM on item descriptions, then aligns item textual embeddings with pretrained ID-based embeddings, and finally merges them into CTR models for enhancement. We selected Robert-base [69] as the PLM as it performs the best in the original paper.
@@
  • ClickPrompt (WWW’24) [11]: ClickPrompt uses a CTR model to encode instances into soft tokens, which are then treated as prefix tokens in a PLM for generating instance-level semantic knowledge. Following the original paper, we selected Robert-base [69] as the PLM in our experiment.
A.4 Backbones
We selected six representative CTR methods as backbones in our experiments, including:
@@ A.4 Backbones
  • FM (ICDM’10) [15]: FM introduces factorization machines to model pairwise feature interactions via factorized parameters.
@@
  • DeepFM (IJCAI’17) [16]: DeepFM combines factorization machines with deep neural networks to jointly learn low-order and high-order feature interactions. We fix the DNN layers at [300, 300, 128] in our experiments.
@@
  • FwFM (WWW’18) [17]: FwFM proposes field-weighted factorization machines that learn field-specific interaction weights to better capture heterogeneous feature relationships.
@@
  • FmFM (WWW’21) [18]: FmFM enhances FM with field-matrix factorization using learnable projection matrices per field pair, enabling more expressive cross-field interactions.
@@
  • FinalMLP (AAAI’23) [19]: FinalMLP designs two Stream Feature Interaction networks to progressively refine feature representations learned by MLP. We select the MLP hidden units size from {400, 500} and the MLP layer number from {2, 3}, respectively.
@@
  • WuKong (ICML’24) [20]: WuKong employs an interaction stack on the WuKong layer to capture feature interactions, where each WuKong layer consists of a Factorization Machine Block (FMB) and a Linear Compress Block. We select the interaction layer number from {4, 8}, the FMB size from {128, 200}, the FMB layer from {2, 3} and the compression dimension from {24, 32} respectively.
@@
The hyperparameter settings of each method are detailed in Appendix A.5. For all the backbone models, we reused the implement of an open-source CTR prediction library [^3] https://github.com/reczoo/FuxiCTR - FuxiCTR [86], following the previous work [19, 87, 47].
A.5 Hyperparameters
A.5.1 Hyperparameter Settings
We use Adam [88] optimizer to train all the CTR models. The learning rate (lr) is set as 0.001 and the weight decay (wd) is tuned in {1e-2, 1e-3, 1e-4, 1e-5, 1e-6, 0}. We fix the feature embedding size to 32 and batch size to 4096 for all the backbones and datasets. For LLaCTR fine-tuning, we fix the batch size as 128, the learning rate of LLM as 1e-4 and rank of LoRA [89] as 8, respectively. We set the feature sampling number as 1000 by default, and set temperature parameter $\tau$ of contrastive loss into 0.02 following previous work [90]. During LLaCTR enhancement, the $\lambda_{kl}$ and $\lambda_{fm}$ are tuned in {0.01, 0.05, 0.1, 0.3, 0.5, 0.7, 1}.
A.5.2 Optimal Hyperparameters
The hyperparameters we searched include: the $\lambda_{kl}$ for Feature Interaction Enhancement (FRE) module, the $\lambda_{fm}$ for Feature Interaction Enhancement (FIE) module, the weight decay (wd) of Adam [88] optimizer and other hyperparameters for backbone CTR models. Table A.2 shows the optimal hyperparameters, including four datasets (Gift Cards, Video Games, Digital Music and MovieLens-1M) and six backbones (FM, DeepFM, FwFM, FmFM, FinalMLP and WuKong).
Table A.2: Optimal hyperparameters for 4 datasets and 6 backbones.
[…]
$\lambda_{fm}$ 0.01 0.01 0.7 0.5
[…]
$\lambda_{fm}$ 0.5 0.5 0.3 0.7
[…]
$\lambda_{fm}$ 0.1 0.1 0.3 0.3
[…]
$\lambda_{fm}$ 0.1 0.1 0.05 0.7
[…]
$\lambda_{fm}$ 0.1 0.3 0.01 0.05
wd 0.01 0.00001 0.000001 0.00001
[…]
MLP layer [2, 3] [2, 2] [2, 3] [2, 3]
[…]
$\lambda_{fm}$ 0.5 0.05 0.5 0.05
wd 0.01 0.00001 0.000001 0.00001
WuKong layer 8 4 8 4
Compr. dim 32 24 32 24
FMB size 200 200 200 200
[…]
Appendix B Supplementary Experiments
B.1 Efficiency Study
Empirical efficiency analysis. Empirically, we analysis the total training time of representative conventional CTR models and LLM-enhanced CTR models on Amazon Video Games and MovieLens-1M datasets with millions of interactions. As shown in Figure B.1, existing LLM-enhanced CTR methods (e.g., KAR, LLM-CF, CTRL and EASE) require over 290 times (on average) more computation time than conventional CTR models. With the increase in the number of interactions, this phenomenon will further exacerbate. In real-world industrial scenarios with interaction scales numbering in the billions, these LLM-enhanced CTR models will face more significant efficiency challenges.
(a) Video Games
(b) MovieLens-1M
Figure B.1: Empirical efficiency study on representative conventional CTR models and LLM-enhanced CTR models. The multiple increase in time cost is reported based on WuKong (the SOTA traditional CTR model).
@@ B.1 Efficiency Study
Efficiency bottleneck analysis. The training time of all LLM-enhanced CTR methods can be roughly divided into four parts: fine-tuning, knowledge generation, semantic encoding and CTR prediction. To study the efficiency bottleneck of the four LLM-enhanced CTR baselines (KAR, LLM-CF, CTRL and EASE) and our LLaCTR, we conducted a detailed training time analysis on the Amazon Video Games and MovieLens-1M datasets.
Table B.3: Running time (s) details of LLM-enhanced CTR baselines and our LLaCTR.
[…]
LLM-CF 15,179 78,104 1,676 178 95,137
CTRL 0 0 5,839 329 6,168
EASE 32,221 0 1,747 125 34,093
[…]
LLM-CF 5,234 26,094 355 64 31,747
CTRL 0 0 1,401 141 1,542
EASE 12,813 0 49 43 12,905
[…]
The results are shown in Table B.3. The mark "/" represents that the corresponding phase does not exist in this method. We can derive:
@@
  • 1) The efficiency bottleneck of KAR and LLM-CF is the knowledge generation. KAR requires to generate knowledge for a large number of users/items, while LLM-CF needs to generate the Chain-of-Thought (CoT) data on numerous instances (even if sampled as 10%). The reasoning and generation of LLMs consume a significant amount of time, and rise constantly as the length of generated text increases.
@@
  • 3) The time overhead of fine-tuning with EASE is higher than other methods. EASE freezes the LLM and trains trains the semantic adaptor network during fine-tuning, but the frozen LLM still needs to be involved in the model forward phase. Besides, the parameter size of EASE’s semantic adaptor network (a BERT-Base [68] network with 12 transformer blocks and an additional 2 cross-attention layers) is about 110M, while the trainable parameter size of LoRA [89] based Llama3-8B [3] (used in LLM-CF and LLaCTR) is only approximately 3.4M (with the rank LoRA as 8), leading to a higher fine-tuning time overhead for EASE.
@@
Time complexity analysis. To further study the training efficiency of LLaCTR and other LLM-enhanced CTR methods, we discuss the training time complexity of these models here. Let $A$ represent the average input token sequence length of the LLM, $B$ represent the average output token sequence length of the LLM, and $D$ represent the embedding dimension of the LLM. The average time overhead of the LLM during fine-tuning and encoding can be expressed as $H(A)$, while the average time overhead during generation can be expressed as $G(A,B)$. For the original CTR model, let $t$ represent the average time overhead of the model and $d$ represent the feature embedding dimension.
@@
We use $U$ to represent the number of users, $I$ to represent the number of items, $N$ to represent the total number of instances, respectively. For LLaCTR, let $K$ represent the number of fields and $S$ represent the feature sample number of SSFT module. The time complexity of the SSFT module is $O(SKH(A)+KH(A))$, while the FRE module is $O(KDd)$ and the FRE module is $O(t+K^{2})$, respectively. The overall time complexity of LLaCTR is $O((S+1)KH(A)+KDd+(t+K^{2}))$. Since the time overhead on the LLM is much greater than the enhancement on CTR models, the SSFT module will dominate the overall time overhead, and the time complexity can be simplified to $O((S+1)KH(A))$. Similarly, the time complexity of KAR can be expressed as $O((U+I)G(A,B)+(U+I)H(A))$, the time complexity of LLM-CF can be expressed as $O(2NH(A)+NG(A,B))$, the time complexity of CTRL can be expressed as $O(NH(A))$, and the time complexity of EASE can be expressed as $O(2IH(A))$. Given that $K<SK\ll(U+I)<N$, LLaCTR exhibits significantly lower training time complexity compared to other LLM-enhanced CTR methods. This further highlights the efficiency of LLaCTR’s field-level enhancement approach.
B.2 Empirical Study
B.2.1 Fine-tuning Loss Study
@@
To select the fine-tuning loss in the SSFT module of our LLaCTR, we compare the performance of language generative loss ($\mathcal{L}_{LG}$) with contrastive loss ($\mathcal{L}_{CL}$) on Amazon Gift Cards and Video Games datasets over DeepFM and WuKong backbones, and the results are shown in Table B.4.
Table B.4: Fine-tuning loss (in SSFT module) comparison between language generative loss ($\mathcal{L}_{LG}$) and contrastive loss ($\mathcal{L}_{CL}$).
[…]
LLaCTR ($\mathcal{L}_{LG}$) 0.4659 0.6794 4.73% 0.5755 0.6718 1.55%
[…]
LLaCTR ($\mathcal{L}_{LG}$) 0.4748 0.6800 4.14% 0.5795 0.6795 1.57%
[…]
We can find that the contrastive loss performs slightly better than the language generative loss, so we use the contrastive loss as the final fine-tuning loss in our experiments.
B.2.2 Alignment Loss Study
@@
To select the feature representation enhancement loss in the FRE module of our LLaCTR, we compare the performance of mean squared error loss ($\mathcal{L}_{MSE}$), contrastive loss ($\mathcal{L}_{CL}$), and Kullback-Leibler divergence loss ($\mathcal{L}_{KL}$) on Amazon Gift Cards and Video Games datasets over DeepFM and WuKong backbones, and the results are shown in Table B.5.
Table B.5: Feature representation enhancement loss (in FRE module) comparison between mean squared error loss ($\mathcal{L}_{MSE}$), contrastive loss ($\mathcal{L}_{CL}$), and Kullback-Leibler divergence loss ($\mathcal{L}_{KL}$).
                    Method Gift Cards         Video Games
[…]
LLaCTR ($\mathcal{L}_{MSE}$) 0.4750 0.6767 3.15% 0.5760 0.6711 1.15%
LLaCTR ($\mathcal{L}_{CL}$) 0.4750 0.6763 2.92% 0.5762 0.6699 0.44%
[…]
LLaCTR ($\mathcal{L}_{MSE}$) 0.4755 0.6737 0.49% 0.5787 0.6795 1.57%
LLaCTR ($\mathcal{L}_{CL}$) 0.4753 0.6738 0.55% 0.5805 0.6788 1.17%
[…]
We can find that the KL divergence loss performs better than the MSE loss and CL loss, so we use the KL divergence loss as the final feature representation enhancement loss in our experiments.
B.3 Ablation Study
We conducted an ablation study on different datasets to study the contributions of each component of LLaCTR. For the Self-supervised Field-feature Fine-tuning (SSFT) module, we removed the entire instructing Fine-tuning component (w/o FT). For Feature Representation Enhancement (FRE) module and Feature Interaction Enhancement (FIE) module, we removed the entire alignment loss (w/o FRE) and the entire field interaction enhancement network (w/o FIE), respectively. The results are presented in Table B.6.
Table B.6: Ablation Study on LLaCTR.
[…]
w/o FIE 0.4667 0.6710 0.5767 0.6652 0.3391 0.7572 0.4019 0.8364
[…]
As can be seen, both the three components are important — removing SSFT, FRE or the FIE module would result in performance drops in most case. Delving deeper into the SSFT module, we observe that self-supervised field-feature fine-tuning on LLMs is indeed helpful. The field description encoded by the fine-tuned LLM has higher quality than that extracted from the pre-trained corpus. For FRE module, we observe that the developed KL-Divergence loss is also important. It can effectively inject semantic knowledge of fields into the feature representations. For FIE module, we find that for both explicit feature modeling and implicit feature modeling CTR methods, injecting field interaction scores can effectively guide feature interaction and bring AUC improvement.
B.4 Hyperparameters Sensitivity Study
Figure B.2 illustrates performance of LLaCTR with different hyper-parameters ($\lambda_{kl},\lambda_{fm}$) on Gift Cards and Video Games datasets, where $\lambda_{kl}$ and $\lambda_{fm}$ control the effects of Feature Representation Enhancement (FRE) and Feature Interaction Enhancement (FIE) respectively.
(a) $\lambda_{kl}$ on Gift Cards
(b) $\lambda_{kl}$ on Gift Cards
(c) $\lambda_{kl}$ on Gift Cards
(d) $\lambda_{fm}$ on Gift Cards
(e) $\lambda_{fm}$ on Gift Cards
(f) $\lambda_{fm}$ on Gift Cards
(g) $\lambda_{kl}$ on Video Games
(h) $\lambda_{kl}$ on Video Games
(i) $\lambda_{kl}$ on Video Games
(j) $\lambda_{fm}$ on Video Games
(k) $\lambda_{fm}$ on Video Games
(l) $\lambda_{fm}$ on Video Games
Figure B.2: Sensitivity analysis w.r.t. $\lambda_{kl}$, $\lambda_{fm}$.

### v2

@@ A.1. Baselines
  • KAR (RecSys’24) (Xi et al., 2024): KAR leverages the LLM as a open-word knowledge base to enhance the profile of users and items, and then pass the encoded user/item embeddings through a knowledge adaptation network to enhance traditional CTR models.
@@
  • LLM-CF (CIKM’24) (Sun et al., 2024): LLM-CF uses LLMs to generate a base of chains of thought (CoT) for sampled instances, which are further retrieved to enhance CTR models. Consistent with the original paper, we sampled 10% of the instances as sampled data for fine-tuning and CoT data generation.
@@
  • CTRL (TORS’23) (Li et al., 2023): CTRL utilizes the LLM as the feature encoder for instances and uses a fine-grained contrastive learning scheme to align cross-modal knowledge.
@@
  • EASE (CIKM’24) (Qiu et al., 2024): EASE trains a semantic adaptor network to align item embeddings with the LLM, and then use it for encoding during inference. Following the original paper, we inited the semantic adapter using BERT-Base (Devlin et al., 2019) with 12 transformer blocks, and cross-attention layers are added to the 6-th and 12-th transformer blocks.
@@
For fair comparisons, unless otherwise specified, we consistently used the Llama3-8B (Dubey et al., 2024) as the LLM in all the LLM-enhanced CTR methods. Besides, we also reproduced two representative PLM-enhanced methods:
@@
  • CELA (arXiv’24) (Wang et al., 2024): CELA first pretrains a PLM on item descriptions, then aligns item textual embeddings with pretrained ID-based embeddings, and finally merges them into CTR models for enhancement. We selected Robert-base (Liu et al., 2019b) as the PLM as it performs the best in the original paper.
@@
  • ClickPrompt (WWW’24) (Lin et al., 2024a): ClickPrompt uses a CTR model to encode instances into soft tokens, which are then treated as prefix tokens in a PLM for generating instance-level semantic knowledge. Following the original paper, we selected Robert-base (Liu et al., 2019b) as the PLM in our experiment.
A.2. Backbones
For all the backbone models, we reused the implement of an open-source CTR prediction library [^3] https://github.com/reczoo/FuxiCTR - FuxiCTR (Zhu et al., 2022), following the previous work (Mao et al., 2023b; Zhu et al., 2023; Li et al., 2024). We selected six representative CTR methods as backbones in our experiments, including:
@@ A.2. Backbones
  • FM (ICDM’10) (Rendle, 2010): FM introduces factorization machines to model pairwise feature interactions via factorized parameters.
@@
  • DeepFM (IJCAI’17) (Guo et al., 2017): DeepFM combines factorization machines with deep neural networks to jointly learn low-order and high-order feature interactions. We fix the DNN layers at [300, 300, 128] in our experiments.
@@
  • FwFM (WWW’18) (Pan et al., 2018): FwFM proposes field-weighted factorization machines that learn field-specific interaction weights to better capture heterogeneous feature relationships.
@@
  • FmFM (WWW’21) (Sun et al., 2021): FmFM enhances FM with field-matrix factorization using learnable projection matrices per field pair, enabling more expressive cross-field interactions.
@@
  • FinalMLP (AAAI’23) (Mao et al., 2023b): FinalMLP designs two Stream Feature Interaction networks to progressively refine feature representations learned by MLP. We select the MLP hidden units size from {400, 500} and the MLP layer number from {2, 3}.
@@
  • WuKong (ICML’24) (Zhang et al., 2024): WuKong employs an interaction stack on the WuKong layer to capture feature interactions, where each WuKong layer consists of a Factorization Machine Block (FMB) and a Linear Compress Block. We select the interaction layer number from {4, 8}, the FMB layer from {2, 3} and the compression dimension from {24, 32} respectively.
@@
Table 4. Fine-tuning loss (in SSFT module) comparison between language generative loss ($\mathcal{L}_{LG}$) and contrastive loss ($\mathcal{L}_{CL}$).
                    Method Gift Cards         Video Games
Logloss ↓ AUC ↑ RelaImpr Logloss ↓ AUC ↑ RelaImpr
DeepFM Base 0.4751 0.6713 0.5762 0.6692
$\mathcal{L}_{LG}$ 0.4659 0.6794 4.73% 0.5755 0.6718 1.55%
$\mathcal{L}_{CL}$ 0.4657 0.6797 4.90% 0.5834 0.6744 3.10%
WuKong Base 0.4755 0.6728 0.5819 0.6767
$\mathcal{L}_{LG}$ 0.4748 0.6800 4.14% 0.5795 0.6795 1.57%
$\mathcal{L}_{CL}$ 0.4745 0.6802 4.24% 0.5763 0.6819 2.91%
Table 5. Feature representation enhancement loss (in FRE module) comparison between mean squared error loss ($\mathcal{L}_{MSE}$), contrastive loss ($\mathcal{L}_{CL}$), and Kullback-Leibler divergence loss ($\mathcal{L}_{KL}$).
                    Method Gift Cards         Video Games
Logloss ↓ AUC ↑ RelaImpr Logloss ↓ AUC ↑ RelaImpr
DeepFM Base 0.4751 0.6713 0.5762 0.6692
$\mathcal{L}_{MSE}$ 0.4750 0.6767 3.15% 0.5760 0.6711 1.15%
$\mathcal{L}_{CL}$ 0.4750 0.6763 2.92% 0.5762 0.6699 0.44%
$\mathcal{L}_{KL}$ 0.4657 0.6797 4.90% 0.5834 0.6744 3.10%
WuKong Base 0.4755 0.6728 0.5819 0.6767
$\mathcal{L}_{MSE}$ 0.4755 0.6737 0.49% 0.5787 0.6795 1.57%
$\mathcal{L}_{CL}$ 0.4753 0.6738 0.55% 0.5805 0.6788 1.17%
$\mathcal{L}_{KL}$ 0.4745 0.6802 4.24% 0.5763 0.6819 2.91%
Appendix B Appendix B: Supplementary Experiments
B.1. Efficiency Study
B.1.1. Efficiency Bottleneck Analysis
@@
The training time of all LLM-enhanced CTR methods can be divided into four parts: fine-tuning, knowledge generation, semantic encoding and CTR prediction. To study the efficiency bottleneck, we conducted a detailed training time analysis on Video Games and MovieLens-1M datasets. The results are shown in Table 6. We can derive:
@@
  • 1) The efficiency bottleneck of KAR and LLM-CF is the knowledge generation. KAR requires to generate knowledge for a large number of users/items, while LLM-CF needs to generate the CoT data on numerous instances (even if sampled as 10%). The reasoning and generation of LLMs consume a significant amount of time, and rise constantly as the length of generated text increases.
@@
  • 3) The time overhead of fine-tuning with EASE is higher than other methods. EASE freezes the LLM and trains trains the semantic adaptor network during fine-tuning, but the frozen LLM still needs to be involved in the model forward phase. Besides, the parameter size of EASE’s semantic adaptor network (a BERT-Base (Devlin et al., 2019) network with 12 transformer blocks and an additional 2 cross-attention layers) is about 110M, while the trainable parameter size of LoRA (Hu et al., 2021) based Llama3-8B (Dubey et al., 2024) (used in LLM-CF and LLaCTR) is only approximately 3.4M (with the rank LoRA as 8), leading to a higher fine-tuning time overhead for EASE.
@@
Table 6. Time Cost (s) details of LLM-enhanced CTR baselines and our LLaCTR.
  Dataset Method        Time Cost in Details
Fine-tuning Generation Enoding Prediction
Video Games KAR 0 77,861 3,886 120
LLM-CF 15,179 78,104 1,676 178
CTRL 0 0 5,839 329
EASE 32,221 0 1,747 125
LLaCTR 599 0 12 124
MovieLens-1M KAR 0 2,057 108 42
LLM-CF 5,234 26,094 355 64
CTRL 0 0 1,401 141
EASE 12,813 0 49 43
LLaCTR 415 0 11 34
B.1.2. Time Complexity Analysis
@@
To further study the training efficiency of LLaCTR and other LLM-enhanced CTR methods, we discuss the training time complexity of these models here. For the LLMs, let $A$ and $B$ denote the average input and output token sequence lengths of the LLM, respectively, and $D$ its embedding dimension. The average time overhead of the LLM can be expressed as $H(A)$ for fine-tuning/encoding and $G(A,B)$ for generation. For the original CTR model, $t$ denotes average time overhead and $d$ the feature embedding dimension. $U$, II, and $N$ denote the number of users, items, and total instances, respectively.
@@
In LLaCTR, let $K$ represent the number of fields and $S$ represent the feature sample number of SSFT module. The time complexity of the SSFT module is $O(SKH(A)+KH(A))$, while the FRE module is $O(KDd)$ and the FRE module is $O(t+K^{2})$, respectively. The overall time complexity of LLaCTR is $O((S+1)KH(A)+KDd+(t+K^{2}))$. Since the time overhead on the LLM is much greater than the enhancement on CTR models, the SSFT module will dominate the overall time overhead, and the time complexity can be simplified to $O((S+1)KH(A))$.
Similarly, the time complexity of KAR is $O((U+I)G(A,B)+(U+I)H(A))$, LLM-CF is $O(2NH(A)+NG(A,B))$, CTRL is $O(NH(A))$, and EASE is $O(2IH(A))$. Given that $K<SK\ll(U+I)<N$, LLaCTR exhibits significantly lower training time complexity compared to other LLM-enhanced CTR methods. This further highlights the efficiency of LLaCTR’s field-level enhancement approach.
B.2. Empirical Study
B.2.1. Fine-tuning Loss Study
@@
For the fine-tuning loss in the SSFT module, we compare the performance of language generative loss ($\mathcal{L}_{LG}$) with contrastive loss ($\mathcal{L}_{CL}$) on over DeepFM and WuKong backbones, and the results are shown in Table 4. We observe that the contrastive loss performs slightly better than the language generative loss, so we use the contrastive loss as the final fine-tuning loss.
Table 7. Ablation Study on the other backbones.
                    Method Gift Games Music Movie
FwFM Base 0.6691 0.6683 0.7553 0.8365
w/o FT 0.6714 0.6757 0.7587 0.8365
w/o FRE 0.6692 0.6729 0.7553 0.8366
w/o FIE 0.6710 0.6652 0.7572 0.8364
LLaCTR 0.6718 0.6780 0.7585 0.8367
FmFM Base 0.6747 0.6706 0.7628 0.8325
w/o FT 0.6759 0.6759 0.7614 0.8334
w/o FRE 0.6750 0.6714 0.7608 0.8334
w/o FIE 0.6753 0.6710 0.7612 0.8332
LLaCTR 0.6761 0.6765 0.7615 0.8335
FinalMLP Base 0.6723 0.6761 0.7557 0.8260
w/o FT 0.6815 0.6787 0.7554 0.8330
w/o FRE 0.6757 0.6772 0.7564 0.8332
w/o FIE 0.6800 0.6775 0.7555 0.8324
LLaCTR 0.6818 0.6832 0.7552 0.8340
B.2.2. Alignment Loss Study
@@
For the feature representation enhancement loss in the FRE module, we compare the performance of mean squared error loss ($\mathcal{L}_{MSE}$), contrastive loss ($\mathcal{L}_{CL}$), and Kullback-Leibler divergence loss ($\mathcal{L}_{KL}$) over DeepFM and WuKong backbones, and the results are shown in Table 5. We find that the KL divergence loss performs better than the MSE loss and CL loss, which is more suitable for enhancement loss.
B.3. Inference Efficiency Study
Table 8. Inference Time (s) on LLM-enhanced CTR methods.
  Dataset WuKong KAR LLM-CF CTRL EASE LLaCTR
Video Games 1.57 1.66 1.84 1.81 1.73 1.63
MovieLens-1M 0.32 0.41 0.53 0.46 0.45 0.41
We compare the total inference time of different LLM-enhanced methods on the Video Games and MovieLens-1M datasets, with the results shown in Table 8 (WuKong serves as the backbone model). It can be observed that the differences among various LLM-enhanced methods during the inference stage are negligible and do not introduce significant latency compared to the backbone model. This is because these LLM-enhanced methods have pre-stored the knowledge of LLMs, requiring only lightweight networks and a small number of parameters to inject the knowledge into the inference process of traditional CTR models, thereby avoiding excessive additional latency.
B.4. Ablation Study
The ablation study results on the FwFM, FmFM and FinalMLP backbones are presented in Table 7. As can be seen, the results are consistent with those obtained from other backbones. Both the three modules are important, removing each would result in performance drops.
B.5. Hyperparameters Sensitivity